A random sample of 128 firms was selected to investigate the average amount spent on liability insurance. The sample mean was found to be £25,500 per annum with a sample standard deviation equal to £3,250.
(i) Calculate the 95% confidence interval for the population mean.
(4 marks)
(ii) Calculate the 99% confidence interval for the population mean.
(4 marks)
(b) Discuss any four non probability sampling methods.
1
Expert's answer
2019-12-10T10:53:10-0500

Here we have "\bar{X}=25500,n=128" and "\sigma=3250"

i) Confidence interval is:

"\bar{X}\pm z_\frac{\alpha}{2}*\frac{\sigma}{\sqrt{n}}"

Here, we have to find "95\%" confidence interval, so "\alpha=0.05"

Now, the value of "z_\frac{\alpha}{2}" Can be calculated using normal distribution table, which comes out to be "1.96"

So, the confidence interval will be

"25500\pm 1.96*\frac{3250}{\sqrt{128}}"

"25500\pm 6370/\sqrt{128}"

"25500\pm 563.033"

"=[24936.967,26063.033]"


ii) Confidence interval is:

"\bar{X}\pm z_\frac{\alpha}{2}*\frac{\sigma}{\sqrt{n}}"

Here, we have to find "99\%" confidence interval, so "\alpha=0.01"

Now, the value of "z_\frac{\alpha}{2}"

Can be calculated using normal distribution table, which comes out to be 2.58

So, the confidence interval will be

"25500\pm 2.58*\frac{3250}{\sqrt{128}}"

"25500\pm 8385/\sqrt{128}"

"25500\pm 741.1363"

"[24758.8637,26241.1363]"


b)

Non-probability sampling represents a group of sampling techniques that help researchers to select units from a population that they are interested in studying. Some of them are:


Quota sampling: With proportional quota sampling, the aim is to end up with a sample where the strata (groups) being studied (e.g., males vs. females students) are proportional to the population being studied.


Convenience sampling: A convenience sample is simply one where the units that are selected for inclusion in the sample are the easiest to access.


Purposive sampling: Purposive sampling, also known as judgmental, selective or subjective sampling, reflects a group of sampling techniques that rely on the judgement of the researcher when it comes to selecting the units.


Self selection sampling: Self-selection sampling is appropriate when we want to allow units or cases, whether individuals or organisations, to choose to take part in research on their own accord.